An astronaut on the moon tosses a ball upwards; when will the ball initially be 35 feet above the lunar surface if the ball's height (feet) follows: h(t)=-2.6t^2+32.1t_5.5, and t is in seconds from release?
a)1 second
b)2 seconds
c)1/2 seconds
d)13.2 seconds
I came up with 1 second and 11.. a little confused.

Answer by stanbon(75887)

An astronaut on the moon tosses a ball upwards; when will the ball initially be 35 feet above the lunar surface if the ball's height (feet) follows:
h(t)=-2.6t^2+32.1t_5.5, and t is in seconds from release?
a)1 second
b)2 seconds
c)1/2 seconds
d)13.2 seconds
-------------------
-2.6t^2 + 32.1t + 5.5 = 35
-2.6t^2 + 32.1t -29.5 = 0
----------
The 1st time the ball reaches 35 ft. (on the way up) is 1 second.
The 2nd time it reaches 35 ft. (on the way down) is 11.35 seconds
==================
Cheers,
Stan H.